If 27litre lemon juice requires 1500ml measured salt how much measured ml salt does 25litres of lemon juice require?
Answer by checkley79(3341)

27/1500=25/X CROSS MULTIPLY.
27X=25*1500
27X=37500
X=37500/27
X=1388.89 ANS.